OpenStudy (anonymous):

Can somebody help me understand and solve this implicit differentiation problem? Find y^1 by implicit differentiation: (x^2)y=2x+siny

OpenStudy (amistre64):

how far do you get?

OpenStudy (amistre64):

and is that really a y^1?

OpenStudy (amistre64):

find y' i think is what it means

OpenStudy (anonymous):

oh no, its derivative, i just wasn't sure how to type it out

OpenStudy (amistre64):

its a single quote :)

OpenStudy (anonymous):

haha thanks

OpenStudy (amistre64):

most of the time, implicits confusion is a result of not having a suitable understanding of a derivative to begin with. no offense

OpenStudy (amistre64):

try this; what is the derivative of x^3 ?

OpenStudy (anonymous):

yea, i admit i have problems with this, lol. i think, 3x^2?

OpenStudy (amistre64):

close, but i never stated the derivative "with respect to x" so you instinctively threw out the derived bit

OpenStudy (amistre64):

it is 3x^2 x'

OpenStudy (amistre64):

try this one, derivative of: 2y^2

OpenStudy (anonymous):

i'll try saying 4y?

OpenStudy (amistre64):

youve got the thought down, but you keep throwing away your derived bit. it is: 4y y'

OpenStudy (anonymous):

oh shoot. yea, i guess i forget that part. how do u get y', again?

OpenStudy (amistre64):

the confusion lies in this: \[\frac{d}{dx}(x^3)=\frac{dx}{dx}\ 3x^2\] but since we derive with respect to "x"; and dx/dx = 1. you are used to discarding it without a thought

OpenStudy (amistre64):

\[\frac{d}{dx}(2y^2)=\frac{dy}{dx}\ 4y\] the derived bit here doesnt go away since you have no idea what its value is

OpenStudy (anonymous):

wow, u understand my train of thought better than i understand myself, lol

OpenStudy (amistre64):

:) as long as you can retrain yourself to do derivatives correctly, implicits pose no problem

OpenStudy (anonymous):

thats true. the thing that sucks about calculus is that it keeps building up on ur previous knowledge :(

OpenStudy (amistre64):

\[x^2y=2x+sin(y)\] \[\frac{d}{dx}[x^2y=2x+sin(y)]\] \[\frac{d}{dx}[x^2y]=\frac{d}{dx}[2x]+\frac{d}{dx}[sin(y)]\] \[\frac{d}{dx}[x^2y]=\frac{dx}{dx}2+\frac{dy}{dx}cos(y)\] \[\frac{d}{dx}[x^2y]=2+\frac{dy}{dx}cos(y)\] the only trick now is how to apply a product rule, but you already know the product rule right?

OpenStudy (amistre64):

[fg]'=f'g+fg' right? what happens when we change the letters? [xy]'=x'y+xy'

OpenStudy (amistre64):

nothing changes

OpenStudy (amistre64):

\[\frac{d}{dx}[x^2y]=2+\frac{dy}{dx}cos(y)\] \[\frac{d}{dx}[x^2]y+x^2\frac{d}{dx}[y]=2+\frac{dy}{dx}cos(y)\] \[\frac{dx}{dx}2xy+x^2\frac{dy}{dx}1=2+\frac{dy}{dx}cos(y)\] \[2xy+x^2\frac{dy}{dx}=2+\frac{dy}{dx}cos(y)\] do you follow that?

OpenStudy (amistre64):

the rest is just algebra to get to the dy/dx

OpenStudy (anonymous):

oh so, would y'=(2-2xy)/x^2-cosy?

OpenStudy (amistre64):

i believe so

OpenStudy (anonymous):

yay! would u mind helping through an even more complex but simlar problem?

OpenStudy (amistre64):

i got the time

OpenStudy (anonymous):

thanks. here it is: find y" by implicit differentiation: x^4y^4=3

OpenStudy (anonymous):

whoops, i meant to put x^4 + y^4 = 3

OpenStudy (amistre64):

\[[x^4+y^4=3]'\] \[[x^4]'+[y^4]'=[3]'\] what would we do from here?

OpenStudy (anonymous):

uh, taking derivative of each: 4x^3(x') + 4y^3(y')=1? I have a feeling its wrong...

OpenStudy (amistre64):

its mostly right :) [3]'=0 since its just a constant

OpenStudy (amistre64):

now solve it so y' is to one side

OpenStudy (anonymous):

okay: y'=4x^3(x')/4y^3?

OpenStudy (amistre64):

watch the algebra, simple mistakes will cost you on a test. and lets go ahead and consider x' to be 1, i just like to keep it there as a reminder of what happens while deriving. y' = -4x^3/4y^3 = -x^3/y^3

OpenStudy (anonymous):

oh i see

OpenStudy (amistre64):

now if your goal is to have the 2nd derivative; then this is just a product rule

OpenStudy (amistre64):

or a quotient rule; your pick :)

OpenStudy (amistre64):

\[y'=-\frac{x^3}{y^3}\]or\[y'=-\left(\frac{x}{y}\right)^3\]or\[y'=-x^3y^{-3}\] we have a few options to choose from and which ever one is easiest for you is the best

OpenStudy (anonymous):

lets go with the first one :)

OpenStudy (amistre64):

very well :) \[[y'=-\frac{x^3}{y^3}]'\] \[[y']'=-[\frac{x^3}{y^3}]'\] \[y''=-\frac{y^3[x^3]'-[y^3]'x^3}{y^{3^2}}\] \[y''=-\frac{y^3[x^3]'-[y^3]'x^3}{y^6}\] how would we complete this?

OpenStudy (anonymous):

sorry, i'm a little confused on the third step

OpenStudy (amistre64):

the 3rd step is just the application of the quotient rule: \[[\frac{t}{b}]'=\frac{bt'-b't}{b^2}\]

OpenStudy (anonymous):

oh... gotcha

OpenStudy (anonymous):

(y(3x^2)-(3y^2)x^3)/y^6?

OpenStudy (amistre64):

needs some practice to it, but heres what I get: \[y''=-\frac{y^3[x^3]'-[y^3]'x^3}{y^6}\] \[y''=-\frac{y^3\ 3x^2\ x'-3y^2\ y'\ x^3}{y^6}\] \[y''=-\frac{3x^2y^3-3x^3y^2\ y'}{y^6}\] and since y' = -x^3/y^3 we can fil that in \[y''=-\frac{3x^2y^3-3x^3y^2\ \frac{-x^3}{y^3}}{y^6}\] \[y''=-\frac{3x^2y^3-3x^3\ \frac{-x^3}{y}}{y^6}\] \[y''=-\frac{3x^2y^3+\frac{3x^6}{y}}{y^6}\] \[y''=-\frac{\frac{3x^2y^4}{y}+\frac{3x^6}{y}}{y^6}\] \[y''=-\frac{\frac{3x^2y^4+3x^6}{y}}{y^6}\] \[y''=-\frac{3x^2y^4+3x^6}{y\ y^6}\] \[y''=-\frac{3x^2y^4+3x^6}{y^7}\] and hopefully i kept track of my signs

OpenStudy (amistre64):

might simplify further to: \[y''=-\frac{3x^2}{y^7}(-y^4-x^4)\] or some such contrivial mess

OpenStudy (amistre64):

that last simplification I did mess the signs on :/ lol

OpenStudy (anonymous):

wow that was a really difficult problem